What is the common difference?
0,-2,-4,-6

Answer by stanbon(75887)

What is the common difference?
0,-2,-4,-6
------------------
To find the common difference of any arithmetic sequence
Pick any number and subtract the number before it.
-4--2 = -4+2 = -2
-6--4 = -6+4 = -2
etc.
Cheers,
Stan H.